OpenStudy (mtalhahassan2):

Differentiate each function. Do not expand an expression before differentiating. a) y=sqrt x^2-3

OpenStudy (frank0520):

use the chain rule \[\frac{ d }{ dx }(\sqrt{x^2 -3})\]

OpenStudy (mtalhahassan2):

ok so then it be like =(x-3^1/2) if we finish the sqrt root

OpenStudy (mtalhahassan2):

@frank0520 are you there??

OpenStudy (frank0520):

No, that is not correct

OpenStudy (mtalhahassan2):

why not

OpenStudy (mtalhahassan2):

x^2/2 -3^1/2

OpenStudy (frank0520):

\[\frac{ d }{ dx }(\sqrt{x^2 -3})\] \[\frac{ d }{ dx }(x^2 -3)^{1/2}\] \[\frac{ 1 }{ 2 }(x^2 -3)^{-1/2} \frac{ d }{ dx }(x^2 -3)\] \[\frac{ 1 }{ 2 }(x^2 -3)^{-1/2} (2x)\] \[x(x^2 -3)^{-1/2}\] \[\frac{ x }{ \sqrt{x^2 -3} }\]

OpenStudy (mtalhahassan2):

can someone help me with it: f(x)=1/(x^2-16)^5

OpenStudy (frank0520):

Try it yourself, this is the same as the first problem. Use the chain rule.

OpenStudy (mtalhahassan2):

ok let me try then

OpenStudy (mtalhahassan2):

i get =1/10(x^2-16)^4